OpenStudy (anonymous):

for n a natural number prove that if n greater than or equal to 2, then n ^3 - n is always divisible by 3

OpenStudy (anonymous):

pls help

OpenStudy (anonymous):

n^3-n =n(n^2-1) =(n-1)* n *(n+1) For any natural no ,n one of n, or (n-1), or (n+1) will always be divisible by 3. So n^3-n will always be divisible 3 for any natural no ,n